# Range Retrieval with Graph-Based Indices

### v2

@@ Abstract.
Retrieving points based on proximity in a high-dimensional vector space is a crucial step in information retrieval applications. The approximate nearest neighbor search (ANNS) problem, which identifies the nearest neighbors for a query (approximately, since exactly is hard), has been extensively studied in recent years. However, comparatively little attention has been paid to the related problem of finding all points within a given distance of a query, the range retrieval problem, despite its applications in areas such as duplicate detection, plagiarism checking, and facial recognition. In this paper, we present a set of algorithms for range retrieval on graph-based vector indices, which are known to achieve excellent performance on ANNS queries. Since a range query may have anywhere from no matching results to thousands of matching results in the database, we introduce a set of range retrieval algorithms based on modifications of the standard graph search that adapt to terminate quickly on queries in the former group, and to put more resources into finding results for the latter group. Due to the lack of existing benchmarks for range retrieval, we also undertake a comprehensive study of range characteristics of existing embedding datasets, and select a suitable range retrieval radius for eight existing datasets with up to 100 million points in addition to the one existing benchmark. We test our algorithms on these datasets, and find up to 100x improvement in query throughput over a naive baseline approach, with 5-10x improvement on average, and strong performance up to 100 million data points.
@@ 1. Introduction
Similarity search within databases of high-dimensional vectors has become increasingly important over the last decade due to the rise of semantic embeddings generated by neural networks and large language models (LLMs). Similarity search is a foundational building block of applications such as search, recommendations, advertising, and retrieval augmented generation (RAG). Correspondingly, there has been an explosion of work on efficient algorithms for approximate similarity search (Jegou et al., 2010; Malkov and Yashunin, 2018; Subramanya et al., 2019; Fu et al., 2019), as exact similarity search is prohibitively expensive due to the curse of dimensionality.
@@
The object of approximate similarity search, or approximate nearest neighbor search (ANNS) is finding the top- most similar embeddings. These top- results can then be post-processed for the desired application. However, there are some applications for which retrieving top- embeddings for a fixed is a poor fit. Applications such as duplicate detection, plagiarism checking, and facial recognition require instead to retrieve all results within a certain radius of a query rather than top-, with some post-processing after to verify whether a match exists out of the retrieved items (Douze et al., 2021; Schroff et al., 2015; MetaAI, 2020). Furthermore, in real-world search applications some queries may have tens of thousands of matches while others have none or very few. In these cases, search with a small radius, or range retrieval, may be used to both differentiate these query types (Szilvasy et al., 2024). Range search can also be a useful subroutine in applications using nearest neighbor graphs, such as clustering or graph learning (Halcrow et al., 2020; Li et al., 2020a).
@@
Range search differs from top- search by the diversity in size of the ground truth solutions for a set of queries. In practice, solutions for a set of queries follow a skewed, Pareto-like distribution where the majority of queries tend to have no results within the chosen radius, while a smaller fraction have a small number of results, and a few outliers have thousands to tens of thousands of results. Queries in the middle category are well served by existing similarity search algorithms, which are already highly optimized and efficient. Thus, a good range search algorithm will match the efficiency of top- search on the middle category of queries, while quickly terminating on queries with no results, and efficiently finding all results for those few queries with thousands of results.
@@
Despite the many applications of range search, there has been very little work on designing algorithms specifically for the task of high-dimensional range search. From a practical perspective, it would be ideal of data structures for top- search could be reused or lightly adapted for range search, enabling an existing data structure to serve both types of queries. Data structures for top- search typically fall into one of two categories—partition-based indices, which partition the dataset into cells and exhaustively search a small number of cells at query time, and graph-based indices, which construct a proximity graph over the data points and use a variant of greedy search to answer queries. Graph-based indices are widely acknowledged as achieving equal or better similarity search performance than partition-based indices (Wang et al., 2021), but almost no work has studied the question of adapting graph-based indices for range retrieval. Furthermore, top- search with IVF indices naturally explores thousands of candidates per query by exhaustively checking the distances between the query point and all the points in each cell, while the number of nodes a graph-based search typically explores is only a small multiple of . It is thus a much less trivial algorithmic task to adapt graph-based search to serve such queries (as we will show in Section 4, naive adaptations of existing search methods are not adequate for the task). The graph structure also suggests potential to terminate queries with no results after examining just a handful of vertices. It is natural then to ask whether graph-based indices can be adapted to efficiently serve range queries.
@@
In this paper, we present a set of algorithms designed for approximate high-dimensional range search on graph-based ANNS indices. Our techniques are modifications of the standard graph search algorithm, meaning that we enable one data structure to efficiently serve both top- queries and range queries. We present techniques that allow a range query to quickly terminate when the query has no results, and to efficiently answer queries with thousands of results within their radius. We devise an early-stopping heuristic that is capable of predicting whether a query has no results after only a few hops in the graph search. For queries with many results, we design two algorithms, doubling search and greedy search, that extend the search path to return more results while minimizing wasted work, and compare and contrast situations where each algorithm dominates the other.
For our experimental evaluation, since there is only one publicly released dataset specifically for range search, we evaluate eight state-of-the-art publicly available metric embedding datasets with up to 100 million embeddings and select a suitable radius for each one. We also perform additional analysis of the characteristics of different range search datasets, and how they connect to the magnitude of improvement our algorithms achieve.
We evaluate our algorithms on the set of nine datasets (eight contributed by us, and one public benchmark), and find that our range retrieval algorithms are capable of up to 100x speedup over naive adaptations of top- search, and in most cases 5-10x speedup. We additionally find that the speedup and scalability of our algorithms extends up to datasets with 100 million points. Figure 1 shows a preview of our experimental results on three embedding datasets of sizes one million to 100 million. In each case, for a fixed average precision, we find at least a 10x speedup in throughput; additionally, our algorithms extend to significantly higher accuracy than the baseline.
@@
In Section 1.1 we cover related work. In Section 2 we cover some needed preliminaries. In Section 3 we explore the characteristics of range search datasets and provide heuristics for choosing an acceptable radius, which we use to adapt eight public embedding datasets for range searching. In Section 4 we present our algorithms for range search on graph-based indices, and in Section 5 we present experimental results.
@@ 1.1. Related Work
Some earlier work addresses range retrieval in high dimensions from a theoretical standpoint (Chazelle et al., 2008), or with only minimal experiments that do not extend to the large embeddings used today (Wang et al., 2013). Theory results on nearest neighbor search using locality sensitive hashing (LSH) also implicitly apply to range search, since their approximation guarantees are usually in the form of finding neighbors within a radius of the distance to the true top- result (Indyk and Motwani, 1998).
On the practical side, Meta AI released a range retrieval dataset aimed towards detecting misinformation (MetaAI, 2020), which was used as one of the competition datasets in the NeurIPS 2021 Big ANN Benchmarks Challenge (Simhadri et al., 2021a), but competitors solved this task using naive adaptations of top- search rather than novel range searching algorithms. A recent work by Szilvasy, Mazare, and Douze (Szilvasy et al., 2024) addresses the combined task of retrieving range search results with one metric and then re-ranking them using a more sophisticated model. They address the question of designing a range search metric (RSM) specifically for bulk search—that is, range search over a group of queries with a fixed computational budget—and benchmark its results on an image search application. They use an inverted file (IVF) index, a common data structure for similarity search (Douze et al., 2024), to serve their database queries.
The Starling system (Wang et al., 2024) is an SSD-resident graph-based ANNS index optimized for I/O efficiency. Their work includes benchmarks on range search metrics, and they use an algorithm similar to the doubling beam search algorithm we present in Section 4. However, the focus of their paper is not on comprehensive benchmarking of range search, and since their implementation is a disk-resident index and ours is fully in-memory, our experiments are not directly comparable to theirs.
@@
Some variants on the high-level ideas behind our range search algorithms—that is, early termination of queries with few results, and extensions of beam search for queries with many results—have been used before in the context of nearest neighbor search. A work by Li et al. (Li et al., 2020b) uses a machine learning model to predict when a query can terminate early and still maintain high accuracy. Another work (Xu et al., 2021) uses graph search with two phases (essentially, an initial cheaper phase and a later, more compute intensive phase for selected queries to gather additional candidates).
@@ 2. Preliminaries
Here, we define the range search problem and the metric for measuring the accuracy of the search. Range searching, conceptually, can be defined as retrieving results within a ball of radius for a given metric space. We then define average precision, which indicates how well a search algorithm finds the correct result given a set of queries.
@@
Commonly-used distance functions include Euclidean distance ( norm), and maximum inner product search, which uses negative inner product as a distance function..
@@
Let be a set of points in -dimensions and a query point set. For , let be the true range neighbors of in . Let be an output of a range search algorithm. Then Average Precision is defined as
@@ 3. Characteristics of Range Search Datasets
In this section, we take eight existing embedding datasets from top- retrieval benchmarks and show how to use them for range search—namely, by finding a radius which yields a suitable distribution. We evaluate the characteristics of each dataset; alongside, we also evaluate the one existing range search benchmark.
For a dataset to be meaningful as a range search benchmark, the same distance must be indicative of a match for each query point. While this is not necessarily the case for all embedding datasets, it is not directly observable without full access to the source material, and even with that access would be incredibly costly to determine. Thus, we use the characteristics of existing benchmarks to determine whether a dataset is suitable, and select a radius if so. Real-world query sets tend to follow a power-law distribution, where most points have no matches, while a handful of points have thousands of matches (Szilvasy et al., 2024; MetaAI, 2020). To be useful as a range search benchmark, a query set should roughly follow this distribution in the number of matches for each query.
Another important aspect of suitability for range search is whether there exists a choice of radius that is “robust” in the sense that small perturbations of the radius do not result in a wildly different match distribution. Our later experiments show that datasets that are more robust by this measure tend to perform better over the baseline measures than ones which are more easily perturbed.
See Figure 2 for a description of each dataset used in our benchmarks.
@@
Source | ||||
| BIGANN | L2 | 128 (uint8) | 10000 | 10000 | SIFT descriptors of images (Jegou et al., 2010) | ||||
| DEEP | L2 | 96 (float) | 10000 | .02 | Image embeddings from the last fully-connected layer of the GoogLeNet model (Baranchuk and Babenko, 2021) | ||||
| MSTuring | L2 | 100 (float) | 100000 | .3 | Bing queries encoded by Turing AGI v5 (Zhang et al., 2019) | ||||
| GIST | L2 | 968 (float) | 10000 | .5 | GIST descriptors of images (Jegou et al., 2010) | ||||
| SSNPP | L2 | 200 (uint8) | 100000 | 96237 | SimSearchNet++ encodings of images (MetaAI, 2020) | ||||
| OpenAI | L2 | 1536 (float) | 10000 | .2 | ArXiv articles embedded using the OpenAI text-embedding-ada-002 model (Greene et al., 2022) | ||||
| Text2Image | IP | 200 (float) | 100000 | -.6 | Image embeddings produced by the Se-ResNext-101 model (Baranchuk and Babenko, 2021) | ||||
| Wikipedia | IP | 768 (float) | 5000 | -10.5 | Wikipedia articles embedded from the cohere.ai multilingual 22-12 model (Reimers, 2022) | ||||
| MSMARCO | IP | 768 (float) | 9374 | -62 | Clicked document-query pairs from the ClueWeb22 website corpus (Chen et al., 2024) |
@@ 3.1. Density of Matches by Dataset
This information is presented for each dataset in Figure 3, which shows Euclidean and inner product datasets separately and plots the percent captured from the largest radius achieving 0% capture to the smallest radius achieving 100% capture. As shown in the figure, there is wide variation in each dataset’s response to perturbations in the radius. Datasets BIGANN, GIST, DEEP, Wikipedia, and MSMARCO are the most robust to perturbations in the interesting range (for million size datasets, the applicable range is around to , as most query points will have no results). In these datasets, it is thus less likely that there will be significant numbers of points near the boundary of the ball of radius , and thus less work to distinguish points that are just inside the boundary from those which are just outside the boundary. We will see this insight at work in the experimental results in Section 5.
Using the insights from this experiment, we choose an appropriate radius for each dataset. The real (non-normalized) choices for each radius are shown in Figure 2.
@@ 3.2. Frequency Distribution of Matches
Given a choice of radius for each dataset, we now evaluate the frequency distribution of the number of range results for each query point. The distributions are found in Figure 4. We observe a general pattern of most queries having no results and a few large outliers, which is more pronounced in BIGANN, DEEP, MSTuring, and Text2Image, and less pronounced in SSNPP, OpenAI, Wikipedia, and MSMARCO. GIST is unusual among the datasets in having many extremely large outliers. For BIGANN, SSNPP, and Wikipedia, we also show the frequency distribution for a larger version of the same base dataset. Since these are effectively larger samples from the same distribution, we observe an increase in density of range results.
| Dataset | ||||||
| BIGANN-1M | 9728 | 143 | 84 | 45 | 0 | 0 |
| BIGANN-10M | 9590 | 173 | 99 | 93 | 45 | 0 |
| BIGANN-100M | 9413 | 212 | 136 | 101 | 93 | 45 |
| DEEP-1M | 9923 | 56 | 19 | 2 | 0 | 0 |
| MSTuring-1M | 95716 | 2443 | 20 | 21 | 0 | 0 |
| GIST-1M | 8487 | 830 | 160 | 143 | 134 | 246 |
| SSNPP-1M | 97422 | 2424 | 254 | 0 | 0 | 0 |
| SSNPP-10M | 91575 | 7310 | 954 | 161 | 0 | 0 |
| SSNPP-100M | 80795 | 13719 | 4357 | 971 | 158 | 0 |
| OpenAI-1M | 7030 | 2564 | 372 | 34 | 0 | 0 |
| Text2Image-1M | 99327 | 669 | 4 | 0 | 0 | 0 |
| Wikipedia-1M | 4445 | 482 | 73 | 0 | 0 | 0 |
| Wikipedia-10M | 3385 | 1328 | 277 | 10 | 0 | 0 |
| MSMARCO-1M | 7022 | 2199 | 152 | 3 | 0 | 0 |
4. Algorithms for Range Search on Graph-Based ANNS Indices

### v3

@@ Abstract.
Retrieving points based on proximity in a high-dimensional vector space is a crucial step in information retrieval applications. The approximate nearest neighbor search (ANNS) problem, which identifies the nearest neighbors for a query, has been extensively studied in recent years. However, comparatively little attention has been paid to the related problem of finding all points within a given distance of a query, the range retrieval problem, despite its applications in areas such as duplicate detection, plagiarism checking, and facial recognition. In this paper, we present new techniques for range retrieval on graph-based vector indices, which are known to achieve excellent performance on ANNS queries. Since a range query may have anywhere from no matching results to thousands of matching results in the database, we introduce a set of range retrieval algorithms based on modifications of the standard graph search that adapt to terminate quickly on queries in the former group, and to put more resources into finding results for the latter group. Due to the lack of existing benchmarks for range retrieval, we also undertake a comprehensive study of range characteristics of existing embedding datasets, and select a suitable range retrieval radius for eight existing datasets with up to 1 billion points in addition to one existing benchmark. We test our algorithms on these datasets, and find up to 100x improvement in query throughput over a standard graph search and the FAISS-IVF range search algorithm. We also find up to 10x improvement over a previously suggested modification of the standard beam search, and strong performance up to 1 billion data points.
@@ 1. Introduction
Similarity search within databases of high-dimensional vectors has become increasingly important over the last decade due to the rise of semantic embeddings generated by neural networks and large language models (LLMs). Similarity search is a foundational building block of applications such as search, recommendations, advertising, and retrieval augmented generation (RAG). Correspondingly, there has been an explosion of work on efficient algorithms for approximate similarity search (Jegou et al., 2010; Malkov and Yashunin, 2018; Subramanya et al., 2019), as exact similarity search is prohibitively expensive due to the curse of dimensionality.
@@
The object of approximate similarity search, or approximate nearest neighbor search (ANNS) is finding the top- most similar embeddings. These top- results can then be post-processed for the desired application. However, there are some applications for which retrieving top- embeddings for a fixed is a poor fit. Applications such as duplicate detection, plagiarism checking, and facial recognition require instead to retrieve all results within a certain radius of a query, with some post-processing after to verify whether a match exists out of the retrieved items (Douze et al., 2021; Schroff et al., 2015; MetaAI, 2020). Furthermore, in real-world search applications some queries may have tens of thousands of matches while others have none or very few. In these cases, search with a small radius, or range retrieval, may be used to both differentiate these query types (Szilvasy et al., 2024). Range search can also be a useful subroutine in applications using nearest neighbor graphs, such as clustering or graph learning (Halcrow et al., 2020; Li et al., 2020a).
@@
Range search differs from top- search by the diversity in size of the ground truth solutions for a set of queries. In practice, ground truth for a set of queries follows a skewed, Pareto-like distribution where the majority of queries tend to have no results within the chosen radius, while a smaller fraction have a small number of results, and a few outliers have thousands to tens of thousands of results. Queries in the middle category are well served by existing similarity search algorithms, which are already highly optimized and efficient. Thus, a good range search algorithm will match the efficiency of top- search on the middle category of queries, while quickly terminating on queries with no results, and efficiently finding results for those few queries with thousands of results.
@@
Despite the many applications of range search, there is very little work on designing algorithms specifically for the task of high-dimensional range search. From a practical perspective, it would be ideal if indices for top- search could be reused or lightly adapted for range search, enabling an existing index to serve both types of queries. Indices for top- search typically fall into one of two categories—partition-based indices, which partition the data set into cells and exhaustively search a small number of cells at query time (Jegou et al., 2010; Indyk and Motwani, 1998), and graph-based indices, which construct a proximity graph over the data points and use a variant of greedy search to answer queries (Fu et al., 2019; Subramanya et al., 2019; Malkov and Yashunin, 2018; Wang et al., 2024). Graph-based indices are widely acknowledged as achieving equal or better similarity search performance than partition-based indices (Wang et al., 2021). However, there is almost no work has studied the question of adapting graph-based indices for range retrieval. Furthermore, top- search with IVF indices naturally explores thousands of candidates per query by exhaustively checking the distances between the query point and all the points in each cell, while the number of nodes a graph-based search typically explores is only a small multiple of . It is thus a much less trivial algorithmic task to effectively adapt graph-based search to serve such queries (as we will show in Section 4, naive adaptations of existing search methods are not adequate for the task). The graph structure also suggests potential to terminate queries with no results after examining just a handful of vertices. It is natural then to ask whether graph-based indices can be adapted to efficiently serve range queries.
@@
In this paper, we present a set of algorithms designed for approximate high-dimensional range search on graph-based ANNS indices. Our techniques are modifications of the standard graph search algorithm, meaning that we enable one data structure to efficiently serve both top- queries and range queries. We devise an early-stopping heuristic that is capable of predicting whether a query has no results after only a few hops in the graph search. For queries with many results, we design an algorithm that extends the search path to return more results while minimizing wasted work, and compare and contrast situations where each algorithm dominates the other.
For our experimental evaluation, since there is only one publicly released dataset specifically for range search, we evaluate eight state-of-the-art publicly available metric embedding datasets with up to 1 billion points and select a suitable radius for each one. We also perform additional analysis of the characteristics of different range search datasets.
We evaluate our algorithms on the set of nine datasets (eight contributed by us, and one public benchmark) against both a simple adaptation of top- search, an in-memory variant of the range search algorithm used in the Starling system (Wang et al., 2024) which we refer to as doubling search, and the range search algorithm from FAISS-IVF. We find that our range retrieval algorithms are capable of up to 100x speedup over the top- and FAISS-IVF baselines, and up to 10x speedup over the doubling search. We additionally find that the speedup and scalability of our algorithms extends up to datasets with 1 billion points. Figure 1 shows a preview of our experimental results on three embedding datasets of sizes one million to 1 billion points.
@@
Sections 1.1 and 2 cover related work and preliminaries, respectively. In Section 3 we explore the characteristics of range search datasets and provide heuristics for choosing an acceptable radius, which we use to adapt eight public embedding datasets for range searching. In Section 4 we present our algorithms for range search on graph-based indices, and in Section 5 we present experimental results.
@@ 1.1. Related Work
Some earlier work addresses range retrieval in high dimensions from a theoretical standpoint (Chazelle et al., 2008), or with only minimal experiments that do not extend to the large embeddings used today (Wang et al., 2013). Theory results on nearest neighbor search using locality sensitive hashing (LSH) also apply to range search, since their approximation guarantees are of the form of finding neighbors within a radius of the distance to the true top- result (Indyk and Motwani, 1998).
Meta AI released a range retrieval dataset aimed towards detecting misinformation (MetaAI, 2020), which was used as one of the competition datasets in the NeurIPS 2021 Big ANN Benchmarks Challenge (Simhadri et al., 2021a), but competitors solved this task using naive adaptations of top- search rather than novel range searching algorithms. A recent work by Szilvasy, Mazare, and Douze (Szilvasy et al., 2024) addresses the combined task of retrieving range search results with one metric and then re-ranking them using a more sophisticated model. They design a range search metric specifically for bulk search—that is, range search over a group of queries with a fixed search budget—and benchmark its results on an image search application. They use an inverted file (IVF) index, a common data structure for similarity search (Douze et al., 2024), to serve their database queries. The well-known FAISS-IVF library includes an option for range search on an inverted file index (Douze et al., 2024).
The Starling system (Wang et al., 2024) is an SSD-resident graph-based ANNS index optimized for I/O efficiency. Their work includes benchmarks on range search metrics, but the focus of their paper is not on comprehensive benchmarking of range search, and since their implementation is a disk-resident index and ours is fully in-memory, our experiments are not directly comparable to theirs. In Section 4 we describe the range search algorithm used in the Starling system, and how we adapted it for the in-memory case and made some small experimental improvements before using it as a baseline.
@@
Some variants on the high-level ideas behind our range search algorithms—that is, early termination of queries with few results, and extensions of beam search for queries with many results—have been used before in the context of nearest neighbor search. Two works (Li et al., 2020b; cha, [n. d.]) use machine learning models to predict when a query can terminate early and still maintain high accuracy. Another work (Xu et al., 2021) uses graph search with two phases (essentially, an initial cheaper phase and a later, more compute intensive phase to gather additional candidates).
Not to be confused with range search, the range filtering problem addresses vectors with numerical labels such as timestamps, and asks to retrieve all vectors within a certain range on the labels (e.g., finding all vectors added to the database within a certain time range). This problem, which deals with external metadata on the vectors rather than finding all vectors within a ball of radius , has been the subject of some recent work (Zuo et al., 2024; Engels et al., 2024; Xu et al., 2024).
@@ 2. Preliminaries
Commonly-used distance functions include Euclidean distance ( norm), and maximum inner product search, which uses negative inner product as a distance function (Manohar et al., 2024).
@@
Let be a set of points in -dimensions and a query point set. For , let be the true range neighbors of in . Let be an output of a range search algorithm, which is only allowed to consist of results that are truly within radius of the query point. Then Average Precision is defined as
Throughout this work, we will use the notation to refer to the set of directed edges , where is a directed graph.
@@ 3. Characteristics of Range Search Datasets
In this section, due to the fact that there is only one existing high-dimensional range search benchmark, we take eight existing embedding datasets from top- retrieval benchmarks and show how to use them for range search by finding a suitable radius. We evaluate the characteristics of each dataset; alongside, we also evaluate the one existing range search benchmark.
For a dataset to be a meaningful range search benchmark, the same distance must consistently indicate a match for each query point. This is not always true for embedding datasets and is difficult to verify without full access to the source material. Thus, we assess existing benchmarks to determine suitability and select a radius accordingly. Real-world query sets typically follow a power-law distribution, where most points have no matches, but a few have thousands (Szilvasy et al., 2024; MetaAI, 2020). Another key factor is whether a dataset has a “robust” radius—small perturbations should not drastically alter match distribution. Our experiments show that more robust datasets tend to outperform those sensitive to radius changes. See Table 1 for a description of each dataset used in our benchmarks.
@@
Source |
| BIGANN | L2 | 128 (uint8) | 10000 | 10000 | Images (Jegou et al., 2010) | SIFT descriptors | ||||||
| DEEP | L2 | 96 (float) | 10000 | .02 | Images (Baranchuk and Babenko, 2021) |
| MSTuring | L2 | 100 (float) | 100000 | .3 | Bing queries (Zhang et al., 2019) | Turing AGI v5 | ||||||
| GIST | L2 | 968 (float) | 10000 | .5 | Images (Jegou et al., 2010) | GIST descriptors | ||||||
| SSNPP | L2 | 200 (uint8) | 100000 | 96237 | Images (MetaAI, 2020) | SimSearchNet++ | ||||||
| OpenAI | L2 | 1536 (float) | 10000 | .2 |
| Text2Image | IP | 200 (float) | 100000 | -.6 | Images (Baranchuk and Babenko, 2021) | Se-ResNext-101 | ||||||
| Wikipedia | IP | 768 (float) | 5000 | -10.5 |
IP | 768 (float) | 9374 | -62 |
SimANS |
@@ 3.1. Density of Matches by Dataset
Figure 2 presents data for each dataset, separately plotting Euclidean and inner product results from the largest radius with 0% capture to the smallest achieving 100% capture. As shown, dataset responses to radius perturbations vary widely. BIGANN, GIST, DEEP, Wikipedia, and MSMARCOWebSearch are the most robust in the relevant range (around to for million-size datasets, where most query points yield no results). In these datasets, fewer points lie near the boundary of radius , reducing the effort needed to distinguish between those just inside or outside. Experimental results in Section 5 illustrate this effect. Using insights from this analysis, we select an appropriate radius for each dataset, with real (non-normalized) values shown in Figure 1
@@ 3.2. Frequency Distribution of Matches
Given a choice of radius for each dataset, we now evaluate the frequency distribution of the number of range results for each query point. The full distributions are found in Table 2. We observe a general pattern of most queries having no results and a few large outliers, which is more pronounced in BIGANN, DEEP, MSTuring, and Text2Image, and less pronounced in SSNPP, OpenAI, Wikipedia, and MSMARCOWebSearch. GIST is unusual among the datasets in having many extremely large outliers. For BIGANN, SSNPP, and Wikipedia, we also show the frequency distribution for a larger version of the same base dataset. Since these are effectively larger samples from the same distribution, we observe an increase in density of range results.
| Dataset | ||||||||
| BIGANN-1M | 9728 | 143 | 84 | 45 | 0 | 0 | ||
| BIGANN-10M | 9590 | 173 | 99 | 93 | 45 | 0 | ||
| BIGANN-100M | 9413 | 212 | 136 | 101 | 93 | 45 | ||
| DEEP-1M | 9923 | 56 | 19 | 2 | 0 | 0 | ||
| MSTuring-1M | 95716 | 2443 | 20 | 21 | 0 | 0 | ||
| GIST-1M | 8487 | 830 | 160 | 143 | 134 | 246 | ||
| SSNPP-1M | 97422 | 2424 | 254 | 0 | 0 | 0 | ||
| SSNPP-10M | 91575 | 7310 | 954 | 161 | 0 | 0 | ||
| SSNPP-100M | 80795 | 13719 | 4357 | 971 | 158 | 0 | ||
| OpenAI-1M | 7030 | 2564 | 372 | 34 | 0 | 0 | ||
| Text2Image-1M | 99327 | 669 | 4 | 0 | 0 | 0 | ||
| Wikipedia-1M | 4445 | 482 | 73 | 0 | 0 | 0 | ||
| Wikipedia-10M | 3385 | 1328 | 277 | 10 | 0 | 0 | ||
7022 | 2199 | 152 | 3 | 0 | 0 |
4. Range Search on Graph-Based ANNS Indices